OpenStudy (anonymous):

Help Plz? Verify the identity.cot (θ - π/2 )= - tanθ

OpenStudy (anonymous):

You can write : \[\cot(\theta - \frac{\pi}{2}) = \cot(-(\frac{\pi}{2} - \theta))\]

OpenStudy (anonymous):

Now : \[\cot(- \theta) = - \cot(\theta)\]

OpenStudy (anonymous):

So now by using this what have you got??

OpenStudy (anonymous):

idk im kinda lost

OpenStudy (anonymous):

This is not the answer, I am just saying to use that formula there..

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

Where??

OpenStudy (anonymous):

so what do I do next

OpenStudy (anonymous):

Where??

OpenStudy (anonymous):

Then use: \[\cot(\frac{\pi}{2} - \theta) = \tan(\theta)\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now show me all the steps..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

cot ( theta - pi / 2 )= cos ( theta - pi / 2 ) / sin ( theta - pi / 2 ) cos ( theta - pi / 2 )= cos(theta) cos (pi/2) + sin(theta) sin (pi/2) = cos(theta) (0) + sin(theta) (1) = sin(theta) sin ( theta - pi / 2 )= sin(theta) cos(pi/2) - cos(theta) sin(pi/2) = sin(theta) (0) - cos(theta) (1) = -cos(theta) so you get, cot ( theta - pi / 2 )= sin(theta) / -cos(theta) cot ( theta - pi / 2 )= -tan(theta)

hartnn (hartnn):

so u have the solution.

OpenStudy (anonymous):

ok so I have my answer now right

OpenStudy (anonymous):

This one is right....

OpenStudy (anonymous):

ok thanks so should I write this whole thing as my answer or can I just write cot ( theta - pi / 2 )= sin(theta) / -cos(theta) cot ( theta - pi / 2 )= -tan(theta)

hartnn (hartnn):

whole thing.

OpenStudy (anonymous):

ok thanks